OpenStudy (anonymous):

www.mathportal.org/calculators/analytic-geometry/distance-and-midpoint-calculator.php?val1=3&val2=-2&val3=1&val4=-6&rb1=dist How does the square root of 4 + 16 turn into 2 times the square root of 5?

OpenStudy (amistre64):

\[\sqrt{ab}=\sqrt{a}~\sqrt{b}\] \[\sqrt{20} = \sqrt{4*5}\]